OpenStudy (anonymous):

What is the equation of a circle whose center is at the origin and whose radius is 16?

OpenStudy (amistre64):

x^2 + y^2 = r^2; is the standard format since the center is already at origin; id go with x^2+y^2=sqrt(16), or simply: x^2 + y^2 = 4

OpenStudy (amistre64):

16*16 = 256 tho; not 4 ..... ack

OpenStudy (amistre64):

\[x^2 + y^2 = 256\]

OpenStudy (anonymous):

typical trick question. make the radius a perfect square see if you take the square root!

OpenStudy (anonymous):

hook line and sinker i fall for these

OpenStudy (anonymous):

oops i forgot the squared x^2 + y^2 =4

OpenStudy (anonymous):

amistre is right

OpenStudy (amistre64):

lol ... yep

OpenStudy (anonymous):

@safiah don't fall for the trap it is \[(x-h)^2+(y-k)^2=r^2\]

OpenStudy (anonymous):

amsitre has it

OpenStudy (amistre64):

at least eventually im right ;)